# MLRC-Bench: Can Language Agents Solve Machine Learning Research Challenges?

### v1

@@ Abstract
Existing evaluation of large language model (LLM) agents on scientific discovery lacks objective baselines and metrics to assess the viability of their proposed methods. To address this issue, we introduce MLRC-Bench, a benchmark designed to quantify how effectively language agents can tackle challenging Machine Learning (ML) Research Competitions. Our benchmark highlights open research problems that demand novel methodologies, in contrast to recent benchmarks such as OpenAI’s MLE-Bench (Chan et al., 2024) and METR’s RE-Bench (Wijk et al., 2024), which focus on well-established research tasks that are largely solvable through sufficient engineering effort. Unlike prior work, e.g., AI Scientist (Lu et al., 2024b), which evaluates the end-to-end agentic pipeline by using LLM-as-a-judge, MLRC-Bench measures the key steps of proposing and implementing novel research methods and evaluates them with newly proposed rigorous protocol and objective metrics. Our curated suite of 7 competition tasks reveals significant challenges for LLM agents. Even the best-performing tested agent (gemini-exp-1206 under MLAB (Huang et al., 2024a)) closes only 9.3% of the gap between baseline and top human participant scores. Furthermore, our analysis reveals a misalignment between the LLM-judged innovation and their actual performance on cutting-edge ML research problems. MLRC-Bench is a dynamic benchmark, which is designed to continually grow with new ML competitions to encourage rigorous and objective evaluations of AI’s research capabilities.
Large Language Model, Agentic AI, Automated Scientific Discovery
@@ 1 Introduction
Table 1: Comparison between MLRC-Bench and existing work on automated scientific discovery in machine learning with LLM agents. $\sim$” means that some but not all of the tasks in that benchmark require the indicated capability. “Compute Constraints” indicates whether the solution code must adhere to specified runtime and GPU memory limitations, an important aspect ignored by most prior work.
Problem Identification Method Proposal Experiment Design Code Implementation Evaluation Method Evaluation Object Compute Constraints
AI Scientist (Lu et al., 2024b) LLM & Human Judge Paper
Can LLMs Generate Novel Research Ideas? (Si et al., 2024) Human Judge Idea Proposal
DiscoPOP (Lu et al., 2024a) Performance-Based Function-Level Code
MLE-Bench (Chan et al., 2024) $\sim$ Performance-Based File-Level Code
MLAgentBench (Huang et al., 2024a) $\sim$ Performance-Based File-Level Code
MLRC-Bench (Ours) Performance-Based Repository-Level Code
Evaluating large language model (LLM) research agents (Baek et al., 2024; Li et al., 2024b; Lu et al., 2024b) has so far been restricted to one of two directions. The first direction involves tasking the agent with end-to-end scientific discovery—proposing a research idea, writing implementation code, running experiments, and eventually producing a full paper as done by AI Scientist (Lu et al., 2024b). One issue with such evaluation is the lack of reliable baseline method that enables objective evaluation of the proposed approach. The second direction, on the other hand, evaluates the agent’s ability to produce code that solves a Kaggle-style machine learning (ML) engineering competition, skipping idea proposal and paper writing altogether (Huang et al., 2024a; Chan et al., 2024). While evaluation in this case is straightforward, this setup rarely demands genuine research novelty beyond existing methods. Consequently, neither of these setups paints a full picture of whether LLM research agents are able to come up with research ideas that are both novel and effective, which we aim to address here.
@@
Therefore, this paper introduces MLRC-Bench as a benchmark to evaluate the ability of LLM-based research agents to propose and implement novel methods. Drawing on tasks from recent ML conference competitions, MLRC-Bench enables the evaluation of both novelty and effectiveness of research agents’ ideas compared to a reliable baseline method and the top human solution. In particular, it emphasizes objective metrics on tasks such as LLM merging (Tam et al., 2024) and machine unlearning (Triantafillou et al., 2024), closely mirroring ongoing research challenges. Moreover, the challenges in MLRC-Bench can dynamically grow by incorporating new competitions emerging from future ML conferences and workshops.
We curate MLRC-Bench starting with 7 competition tasks. We pick tasks that involve novel and high-impact problems, spanning areas including LLM safety, multimodal perception, and few-shot learning. Our experimental findings reveal that even the best-performing tested LLM agents, such as gemini-exp-1206 (Pichai, 2024) under the MLAB (Huang et al., 2024a) scaffolding, closes only 9.3% of the gap between baseline and top human participant score. Additionally, our analysis highlights a poor correlation between the novelty judged by LLM and practical effectiveness of agents’ solutions, questioning the reliability of LLM-as-a-judge for research idea evaluation. These results underscore the limitations of current AI research agents in generating and implementing innovative ML solutions, providing a crucial benchmark for future advancements.
Our contributions can be summarized as below:
  • We introduce MLRC-Bench, a dynamic benchmark suite curated from ML conference competitions, featuring open research problems that are both impactful and objectively measurable, and that demand the development of novel methodologies.
  • We conduct large-scale, objective evaluations for a wide array of prominent LLMs with representative agent scaffoldings, highlighting their inability to propose and implement innovative solutions with notable performance gains.
  • We quantify the gap in subjective evaluations of LLM-based research agents, by showing that the perceived idea novelty is misaligned with practical effectiveness.
@@
LLM Merging (Tam et al., 2024) NeurIPS 2024 Efficient LLM Text Accuracy, ROUGE 1 hour 48 GB
Backdoor Trigger Recovery (Xiang et al., 2024) NeurIPS 2024 LLM Safety Text REASR, Recall 0.5 hour 48 GB
Temporal Action Localisation (Heyward et al., 2024) ECCV 2024 Workshop Multimodal Perception Video, Audio mAP 0.5 hour 16 GB
Rainfall Prediction (Gruca et al., 2022) NeurIPS 2023 AI for Science Satellite Data Critical Success Index 0.5 hour 48 GB
Machine Unlearning (Triantafillou et al., 2024) NeurIPS 2023 Data Privacy Image Forgetting Quality, Accuracy 0.5 hour 16 GB
Next Product Recommendation (Jin et al., 2023) KDD Cup 2023 Recommendation System Text Mean Reciprocal Rank 0.5 hour 16 GB
Cross-Domain Meta Learning (Carrión-Ojeda et al., 2022) NeurIPS 2022 Few-Shot Learning Image Accuracy 3.5 hours 16 GB
@@ 2 Related Work
Table 1 presents a summary of the differences between our benchmark and existing work on automating ML research workflow with LLM agents. Scientific discovery in machine learning typically includes four main stages: Problem Identification, where gaps in existing methods are recognized; Method Proposal, which introduces a new approach to address the issue; Experiment Design, involving the selection of datasets, baselines, and metrics for evaluation; and Code Implementation, where the method is realized through executable code. While prior work (Lu et al., 2024b; Si et al., 2024) covers Problem Identification and Experiment Design, evaluation could be subjective based on idea proposal or final paper. Instead, MLRC-Bench focuses on the critical stages of proposing and implementing novel methods, enabling objective performance assessment.
@@
While there are recent benchmarks that focus on code generation in machine learning domain, they do not always require methodological innovation. Works like MLAgentBench (Huang et al., 2024a) and MLE-Bench (Chan et al., 2024) evaluate agents on Kaggle-style ML tasks but prioritize code implementation over novel research contributions. Broader benchmarks such as ScienceAgentBench (Chen et al., 2024b) and DiscoveryBench (Majumder et al., 2024) span multiple scientific domains but lack granularity for ML-specific challenges, while CHIME (Gao et al., 2024) and OpenD5 (Du et al., 2024) target auxiliary tasks like literature review or hypothesis generation. DSBench (Jing et al., 2024) and AAAR-1.0 (Lou et al., 2024) extend evaluations to data science and general R&D workflows but still fall short of addressing cutting-edge ML research innovation. Unlike DiscoPOP (Lu et al., 2024a) and DA-Code (Huang et al., 2024b), which focus on function-level coding or data science, MLRC-Bench requires repository-level code comprehension and generation. RE-Bench (Wijk et al., 2024) and MLGym (Nathani et al., 2025) provide collections of ML research task environments, but their problems either fail to represent most recent research directions (e.g., image classification with CIFAR-10 (Krizhevsky et al., 2009)), or only cover a narrow range of research domains. Besides, existing benchmarks often fail to specify computation constraints (e.g. runtime and GPU memory limit), which are important to encourage efficient yet effective solutions. Because many of these benchmarks are static, they risk becoming contaminated soon after the release of next-generation LLMs. In contrast, our proposed benchmark emphasizes verifiable performance gains on impactful, unsolved research problems, and it can be updated by replacing solved challenges with new competitions to track progress against human experts.
@@
Frameworks like The AI Scientist (Lu et al., 2024b) and MLR-Copilot (Li et al., 2024b) automate end-to-end research workflows but rely largely on subjective reviews of papers or research proposals for evaluating success. In parallel, ResearchAgent (Baek et al., 2024) iteratively refines ideas through multi-agent feedback, and Chain-of-Idea-Agent (Li et al., 2024a) organizes literature into progressive chains to stimulate ideation. However, it remains unclear how subjectively evaluated “novel” ideas translate into actual performance gains. In contrast, we explicitly investigate how such subjective assessments of novelty or idea quality align—or fail to align—with measurable performance improvements. By anchoring evaluations in real-world ML conference competitions, MLRC-Bench emphasizes the importance of balancing subjective judgments with concrete, performance-driven benchmarks—thus closing a critical gap in the evaluation landscape for AI research agents.
@@ 3 MLRC-Bench
Figure 1: Overview of our MLRC-Bench and the evaluation pipeline.
@@ 3.1 Task Selection
  • Non-Trivial: The problem must involve complexity so that it will not be solved by simply applying standard ML algorithms, e.g., calling the XGBoost classifier (Chen & Guestrin, 2016) on a new dataset or prompt engineering with LLMs.
@@
The current version of MLRC-Bench comprises 7 tasks adapted from competitions. These tasks represent a diverse landscape of applied ML research. As shown in Table 2, topics range from LLM safety to multimodal perception, ensuring that benchmarks stress multiple facets of algorithmic creativity. Besides, these tasks are cutting-edge. We primarily select competitions released within the past one year, whose problems remain actively researched, preventing trivial or purely off-the-shelf solutions. In addition, we have the following two considerations for MLRC-Bench design.
@@
MLR-Bench is designed to be updatable. Future releases will incorporate new competitions from the latest conferences and retire old ones if model performance saturates. This mechanism helps track agent progress on fresh, unsolved research questions, mirroring the continually advancing state of ML research.
@@
Many of these competitions occurred after the pre-training cut-off dates for major LLMs. As aforementioned, we will also regularly update the benchmark with the latest suitable competitions. As a result, there is a low chance that the top-performing solutions have already been memorized. This helps ensure that the benchmark measures the agent’s genuine research and development skills rather than simple data retrieval and reproduction.
@@
MLR-Bench specifies explicit computation constraints—such as runtime and GPU memory limits—for each task, mirroring real-world competition scenarios. These standardized constraints promote fair comparisons across different agents, preventing any from gaining unfair advantages through excessive computational resources. By enforcing these limitations, we encourage agents to propose and implement efficient yet effective methods, thereby discouraging trivial or brute-force solutions and incentivizing genuine innovation.
@@ 3.2 Task Environment
  • Starter Code. Refactored from official competition repositories, it contains:
    • A simple, baseline model for comparison.
    • A python environment with the necessary ML frameworks/packages.
    • Scripts for training, inference, and offline or online evaluation.
    • Train, development and test data splits. Training data may not be available for some competitions.
@@
Because our primary goal is to focus on method development, we simplify ML experimentation by unifying the code execution pipeline. Some competition starter kits originally feature complex or inconsistent file structures, so we refactor each one into a standardized, well-organized format, comparable to common ML research project layouts (Figure 1). [^1] See a concrete example here. The resulting codebase allows users to launch experiments with a single command: python main.py --method my_awesome_method --phase dev/test, which applies the specified method to the task and evaluates the result in both development and test phases. To ensure a fair comparison and preserve the integrity of evaluations, the repository enforces file permission management: agents may only modify the methods/ directory (where the algorithmic logic resides in MyMethod.py), while evaluation scripts remain read-only. This setup both promotes methodological innovation and prevents tampering. Additionally, files containing held-out test data are invisible to agents during development phase.
@@ 3.3 Objective Evaluation Metrics
MLRC-Bench supports objective evaluation based on model performance, which provides a clear and measurable way to assess agents’ capabilities of approaching ML research challenges. Moreover, we assess the complexity of an agent’s ideas and the efficiency of their implementations. Importantly, we believe our evaluation supports an easy yet effective and reproducible evaluation on agents’ research novelty. For instance, more innovative approaches may involve rich reasoning and sophisticated mechanisms to push the research frontier, thus more complex. On the other hand, efficiency of implementation highlights whether the agents consider the importance of optimizing these ideas for real-world application. By incorporating performance, complexity, and efficiency evaluations, we ensure that the novelty of an idea is not just theoretical but also practical, offering a more objective measure of its value.
Effectiveness.
Effectiveness is paramount because an ineffective method offers little practical value, regardless of its efficiency or simplicity. Therefore, the core metric is a single performance metric (e.g., accuracy) defined by the competition organizer.
Efficiency.
Competitions typically set baseline compute limits, but faster methods are preferred. We measure each solution’s runtime during training (if applicable) and inference.
Simplicity.
Inspired by standard practice in software estimation (Nguyen et al., 2007), we measure simplicity of agent’s solutions in terms of logical lines of code (LLoC). LLoC excludes comments and blank lines, focusing on executable statements. This metric, while imperfect, offers a rough gauge of code complexity and maintainability (Bhatt et al., 2012). For better readability, we refer to LLoC as “lines of code” throughout this paper.
Table 3: For each research competition and agent, we report the test-phase relative improvement to human, defined as the agent’s solution margin over the baseline normalized by the top human solution’s margin over the baseline, taking the best of 8 trials. Top human participants in each competition will score 100.0 due to the normalization. Additionally, we evaluate two other gpt-4o-based pipelines: MLAB augmented with ideas from either CoI-Agent (Li et al., 2024a) or humans. Best performing agent in each task is highlighted in bold. Our results indicate that providing additional ideas, whether sourced from AI or humans, does not consistently yield performance improvements. The best-performing configuration—gemini-exp-1206 under MLAB—achieves only 9.3% of the human-level improvement over baseline on average, underscoring the inherent difficulty of these research tasks. See Table 4 in Appendix B for absolute improvements to baseline.

### v2

@@ Abstract
We introduce MLRC-Bench, a benchmark designed to quantify how effectively language agents can tackle challenging Machine Learning (ML) Research Competitions, with a focus on open research problems that demand novel methodologies. Unlike prior work, e.g., AI Scientist [36], which evaluates the end-to-end agentic pipeline by using LLM-as-a-judge, MLRC-Bench measures the key steps of proposing and implementing novel research methods and evaluates them with rigorous protocol and objective metrics. Our curated suite of 7 competition tasks reveals significant challenges for LLM agents. Even the best-performing tested agent (gemini-exp-1206 under MLAB [20]) closes only 9.3% of the gap between baseline and top human participant scores. Furthermore, our analysis reveals a misalignment between the LLM-judged innovation and their actual performance on cutting-edge ML research problems. MLRC-Bench is a dynamic benchmark, which is designed to continually grow with new ML competitions to encourage rigorous and objective evaluations of AI’s research capabilities. Our leaderboard and code are publicly available at https://huggingface.co/spaces/launch/MLRC_Bench.
@@ 1 Introduction
Evaluating large language model (LLM) research agents [2, 32, 36] has so far been restricted to two directions. One involves tasking the agent with end-to-end scientific discovery—proposing a research idea, writing implementation code, running experiments, and eventually producing a full paper as done by AI Scientist [36]. One issue with such evaluation is the lack of reliable baseline method that enables objective evaluation of the proposed approach. The second direction, on the other hand, evaluates the agent’s ability to produce code that solves a Kaggle-style machine learning (ML) engineering competition, skipping idea proposal and paper writing altogether [20, 7]. While evaluation in this case is straightforward, this setup rarely demands genuine research novelty beyond existing methods. Consequently, neither of these setups paints a full picture of whether LLM research agents can design research ideas that are both novel and effective, which we aim to address here.
@@
Therefore, this paper introduces MLRC-Bench as a benchmark to evaluate the ability of LLM-based research agents to propose and implement novel methods. Drawing on tasks from recent ML conference competitions, MLRC-Bench enables the evaluation of both novelty and effectiveness of research agents’ ideas compared to a reliable baseline method and the top human solution. In particular, it emphasizes objective metrics on tasks such as LLM merging [47] and machine unlearning [48], closely mirroring ongoing research challenges. Moreover, the challenges in MLRC-Bench can dynamically grow by incorporating new competitions from future ML conferences.
@@
LLM Merging [47] NeurIPS 2024 Efficient LLM Text Accuracy, ROUGE 1 hour 48 GB
Backdoor Trigger Recovery [51] NeurIPS 2024 LLM Safety Text REASR, Recall 0.5 hour 48 GB
Temporal Action Localisation [18] ECCV 2024 Workshop Multimodal Perception Video, Audio mAP 0.5 hour 16 GB
Rainfall Prediction [17] NeurIPS 2023 AI for Science Satellite Data Critical Success Index 0.5 hour 48 GB
Machine Unlearning [48] NeurIPS 2023 Data Privacy Image Forgetting Quality, Accuracy 0.5 hour 16 GB
Next Product Recommendation [26] KDD Cup 2023 Recommendation System Text Mean Reciprocal Rank 0.5 hour 16 GB
Cross-Domain Meta Learning [6] NeurIPS 2022 Few-Shot Learning Image Accuracy 3.5 hours 16 GB
We curate MLRC-Bench starting with 7 competition tasks as shown in Table 1. We pick tasks that involve novel and high-impact problems, spanning areas including LLM safety, multimodal perception, and few-shot learning. Our experimental findings reveal that even the best-performing tested LLM agents, such as gemini-exp-1206 [42] under the MLAB [20] scaffolding, closes only 9.3% of the gap between baseline and top human participant score. Additionally, our analysis highlights a poor correlation between the novelty judged by LLM and practical effectiveness of agents’ solutions, questioning the reliability of LLM-as-a-judge for research idea evaluation. These results underscore the limitations of current AI research agents in generating and implementing innovative ML solutions, providing a crucial benchmark for future advancements.
Our contributions can be summarized as below: 1) We introduce MLRC-Bench, a dynamic benchmark suite curated from ML conference competitions, featuring open research problems that are both impactful and objectively measurable, and that demand the development of novel methodologies; 2) We conduct large-scale, objective evaluations for a wide array of frontier LLMs with representative agent scaffoldings, highlighting their inability to propose and implement innovative solutions with notable performance gains; 3) We pinpoint the flaws in subjective evaluations of LLM-based research agents, by showing that the LLM-judged idea novelty is misaligned with empirical effectiveness.
@@ 2 Related Work
While there are recent benchmarks that focus on code generation in machine learning domain, they do not always require methodological innovation. Works like MLAgentBench [20] and MLE-Bench [7] evaluate agents on Kaggle-style ML tasks but prioritize code implementation over novel research contributions. MLE-Bench requires the final submission to be a CSV file, limiting the data modality. Broader benchmarks such as ScienceAgentBench [13] and DiscoveryBench [38] span multiple scientific domains but lack granularity for ML-specific challenges, while CHIME [16] and OpenD5 [14] target auxiliary tasks like literature review or hypothesis generation. DSBench [27] and AAAR-1.0 [34] extend evaluations to data science and general R&D workflows but still fall short of addressing cutting-edge ML research innovation. RE-Bench [50] and MLGym [39] provide collections of ML research task environments, but their problems either fail to represent most recent research directions (e.g., image classification with CIFAR-10 [30]), or only cover a narrow range of research domains. For example, 6 out of 7 tasks in RE-bench are related to language modeling. Besides, existing benchmarks often fail to specify computation constraints (e.g. runtime and GPU memory limit), which are important to encourage efficient yet effective solutions. The differences between our benchmark and existing work on automating ML research workflow with LLM agents are summarized in Table 3 of Appendix B.
Unlike DiscoPOP [35] and DA-Code [21], which focus on function-level coding or data science, MLRC-Bench requires repository-level code comprehension and generation, thus better reflecting the capabilities required to work with realistic and complex research codebases. This design enables substantial project-level coding tasks, such as generating and editing multiple files [37] and reusing existing utility functions [33]. Moreover, repository-level design supports multi-agent collaboration, where specialized agents for tasks like literature survey, idea brainstorming, code writing, and performance evaluation can iteratively work together [19].
@@
Automated end-to-end research workflows like The AI Scientist [36, 52] and MLR-Copilot [32] rely largely on subjective reviews of papers or research proposals for evaluating success. In parallel, ResearchAgent [2] iteratively refines ideas through multi-agent feedback, and Chain-of-Idea-Agent [31] organizes literature into progressive chains to stimulate ideation. However, it remains unclear how subjectively evaluated “novel” ideas translate into actual performance gains. In contrast, we explicitly investigate how such subjective assessments of novelty or idea quality align—or fail to align—with measurable performance improvements.
@@ 3.1 Task Selection
  • Non-Trivial: The problem must involve complexity so that it will not be solved by simply applying standard ML algorithms, e.g., calling the XGBoost classifier [11] on a new dataset or prompt engineering with LLMs.
@@
The current version of MLRC-Bench comprises seven tasks adapted from existing competitions. [^1] While numerous competitions have been hosted at recent ML/AI conferences, only a limited subset was selected due to factors such as certain challenges being considered solved with the rapid advancement of foundation models, missing or irreproducible evaluation data/code, qualitative rather than quantitative evaluation, insufficient recency, or limited emphasis on algorithmic innovation. These tasks represent a diverse landscape of applied ML research. As shown in Table 1, topics range from LLM safety to multimodal perception, ensuring that benchmarks stress multiple facets of algorithmic creativity. Besides, these tasks are cutting-edge. We select competitions released in recent years, whose problems remain actively researched, preventing trivial or purely off-the-shelf solutions. In addition, we have the following two considerations for MLRC-Bench design.
@@
MLRC-Bench is designed to evolve over time. We will continue to expand the benchmark with tasks from recent conference competitions and retire older ones once model performance saturates. This ensures the benchmark remains aligned with the frontier of machine learning research, enabling continuous tracking of agent progress on fresh and unsolved challenges. In parallel, we have released templates and detailed instructions to facilitate community contributions. [^2] https://tinyurl.com/MLRC-Task-Template All competitions follow a uniform and standardized code structure (detailed in Section 3.2) to streamline research and accelerate progress on LLM agents.
@@
Although most academic competitions require top teams to submit solution reports, the actual implementation code may not be publicly available, reducing the risk of such code appearing in the pretraining data of LLMs. As previously discussed, we will regularly update the benchmark with suitable recent competitions, ensuring that it assesses an agent’s genuine research capabilities rather than its capacity for mere data retrieval and replication.
@@
MLRC-Bench specifies explicit computational constraints, such as runtime and GPU memory limits, closely mirroring real-world competition scenarios. These standardized constraints ensure fairness by preventing agents from leveraging excessive computational resources to gain undue advantage, thereby incentivizing efficient and innovative solutions.
Figure 1: Overview of MLRC-Bench and the evaluation pipeline.
@@ 3.2 Task Environment
  • Starter Code. Refactored from official competition repositories, it contains: 1) a simple, baseline model for comparison; 2) a python environment with the necessary ML frameworks/packages; 3) scripts for training, inference, and offline or online evaluation; 4) train, development and test data splits. Training data may not be available for some competitions.
@@
Because our primary goal is to focus on method development, we simplify ML experimentation by refactoring each competition starter kit into a standardized, well-organized format, comparable to common ML research project layouts (Figure 1). The resulting codebase allows users to launch experiments with a single command: python main.py --method my_awesome_method --phase dev/test, which applies the specified method to the task and evaluates the result in both development and test phases. To ensure a fair comparison and preserve the integrity of evaluations, the repository enforces file permission management: agents may only modify the methods/ directory (where the algorithmic logic resides in MyMethod.py), while evaluation scripts remain read-only. Additionally, files containing held-out test data are invisible to agents during development phase.
@@ 3.3 Objective Evaluation Metrics
MLRC-Bench supports objective evaluation based on three dimensions. We measure Effectiveness by the performance metric (e.g., accuracy) defined by the competition organizer, Efficiency by the solution runtime during training (if applicable) and inference, and Simplicity in terms of logical lines of code (LLoC) [40], inspired by standard practice in software estimation [40]. LLoC excludes comments and blank lines, focusing on executable statements. This metric, while imperfect, offers a rough gauge of code complexity and maintainability [3]. For better readability, we refer to LLoC as “lines of code” throughout this paper.
Main Metric: Relative Improvement to Human Solution.
Quantitative performance comparisons across competitions can be tricky, as each task may differ significantly in its intrinsic difficulty, and the official baseline may be weaker or stronger. To address this, we use the Relative Improvement to Human Solution as our main leaderboard metric that convert each raw performance score $s_{\text{agent}}$ into a normalized score $s^{{}^{\prime}}_{\text{agent}}$, using a linear transformation [5, 50]. Therefore, the score of the baseline solution will be 0, and the top human solution in competition is set to 100. Formally, the normalization is computed as:
3.4 Evaluation Protocol
Our evaluation protocol is designed to prevent AI agents from test set overfitting. Agents will submit their implementation in the form of an edited codebase, particularly within their proposed method in the methods/ directory. Specifically, in a single trial, an agent can iteratively modify the codebase multiple times. We store snapshots of the codebase immediately after each change. Whenever an execution occurs on the development set, we record the resulting metrics and the name of evaluated method for that snapshot. At the end of this iterative development phase, we pick the snapshot with the best development performance i.e., Effectiveness. We then evaluate the method contained in that snapshot on the test set for our final result. [^3] Concretely, we execute the command python main.py --method best_dev_method --phase test. This approach strictly follows standard ML practice and ensures reproducible experimentation. Future work may explore more sophisticated multi-objective selection criteria that additionally weigh runtime (Efficiency) or lines of code (Simplicity) of implementations.
Table 2: For each competition and agent, we report the test-phase relative improvement to the human solution. Best performing agent in each task is highlighted in bold. Our results indicate that providing additional ideas, whether sourced from AI or humans, does not consistently yield performance improvements. The best-performing configuration—gemini-exp-1206 under MLAB—achieves only 9.3% of the human-level improvement over baseline on average, underscoring the inherent difficulty of these research tasks. See Table 4 in Appendix D for absolute improvements to baseline solution.